# SCA-LLM: Spectral-Attentive Channel Prediction with Large Language Models in MIMO-OFDM

### v1

@@ Abstract
In recent years, the success of large language models (LLMs) has inspired growing interest in exploring their potential applications in wireless communications, especially for channel prediction tasks. However, directly applying LLMs to channel prediction faces a domain mismatch issue stemming from their text-based pre-training. To mitigate this, the “adapter + LLM” paradigm has emerged, where an adapter is designed to bridge the domain gap between the channel state information (CSI) data and LLMs. While showing initial success, existing adapters may not fully exploit the potential of this paradigm. To address this limitation, this work provides a key insight that learning representations from the spectral components of CSI features can more effectively help bridge the domain gap. Accordingly, we propose a spectral-attentive framework, named SCA-LLM, for channel prediction in multiple-input multiple-output orthogonal frequency division multiplexing (MIMO-OFDM) systems. Specifically, its novel adapter can capture finer spectral details and better adapt the LLM for channel prediction than previous methods. Extensive simulations show that SCA-LLM achieves state-of-the-art prediction performance and strong generalization, yielding up to $2.4~\text{dB}$ normalized mean squared error (NMSE) advantage over the previous LLM based method. Ablation studies further confirm the superiority of SCA-LLM in mitigating domain mismatch.
Index Terms:
channel prediction, large language models (LLMs), domain mismatch, MIMO-OFDM.
@@ I Introduction
Multiple-input multiple-output orthogonal frequency division multiplexing (MIMO-OFDM) [1] is a key technology offering reliable, high-speed wireless communication capabilities in 5G and future communication systems [2, 3]. This approach combines the spatial benefits of MIMO [4, 5] with the ability of OFDM to handle frequency-selective fading [6]. As a result, significant gains in spectral efficiency and link stability are achieved in MIMO-OFDM systems [1]. To further improve performance, adaptive techniques such as dynamic resource allocation [7], precoding [8], and beamforming[9, 10] are often employed in modern deployments. The success of these techniques, depends heavily on the availability of accurate and timely channel state information (CSI) at the transceiver. However, the acquisition of CSI is often challenging in practice due to the rapid variations of wireless channels over time. As a result, the acquired CSI quickly becomes outdated—a phenomenon known as channel aging [11], particularly in high mobility scenarios or high round-trip scenarios such as non-terrestrial networks (NTN). In addition, frequent channel estimation procedures demand substantial pilot overhead consuming valuable communication resources[12]. To address these issues, channel prediction has emerged as a promising technique. By forecasting future channel conditions, proactive optimization of transmission strategies becomes possible. Consequently, the negative effects of channel aging can be mitigated and the need for frequent pilot transmissions can be significantly reduced. This allows adaptive MIMO-OFDM systems to better realize their full potential.
@@ I-A Prior Works
Channel prediction methods can be broadly categorized into traditional methods and deep learning (DL) based methods. Traditional methods are typically based on explicit mathematical models derived from statistical assumptions or simplified representations of the underlying channel physics. However, it is often difficult for these methods to perform well in complex and dynamic wireless environments. For instance, although advanced model-based techniques, such as the Prony-based angle-delay domain (PAD) method [13], are designed to capture specific channel structures and potentially offer improved accuracy under certain conditions, fundamental limitations are often encountered. While analytical guarantees are provided by PAD, frequent updates to its model parameters are required for adaptation to environmental changes[14]. Substantial computational overhead is incurred by this process as well. There is still a need for more robust and generalizable channel prediction approaches capable of effective adaptation to diverse environments.
@@
In response to this need, DL based methods [14, 15, 16, 17, 18, 19, 20] have emerged as a promising alternative. Rather than relying on predefined mathematical models, DL methods are data-driven approaches that learn complex dependencies and predictive patterns directly from historical channel observations. Architectures such as recurrent neural networks (RNNs) [16], long short-term memory (LSTM) networks [18], gated recurrent units (GRUs) [19], and transformers [20] have been employed, attempting to model the complex spatio-temporal dynamics of wireless channels. For example, focusing on the angle-delay domain for high-mobility massive MIMO systems, a complex-valued neural network (CVNN) based prediction method has been introduced in [14]. This data-driven approach performs element-wise prediction of the angle-delay channel response matrix (ADCRM). By learning directly from historical channel data represented in the angle-delay domain, this method achieved improved prediction accuracy compared to the traditional model-based PAD method [13]. Furthermore, a transformer-based method for channel prediction was introduced in [20]. This approach uses the attention mechanism within the transformer model to establish the relationship between past CSI and future channels. It gives more importance to the most relevant past information, which improves the prediction accuracy and overall communication performance. Additionally, a unified pilot-to-precoder scheme has been proposed either in [20], using a single transformer model to integrate channel estimation, prediction, and precoding, thereby reducing computational complexity. While these approaches have often demonstrated improved accuracy over traditional methods, they still have limitations in practice. Specifically, these models are typically trained for specific tasks and environments. When encountering unseen communication environments without significant retraining, they often exhibit a lack of robustness due to limited sequence modeling capabilities. Therefore, there is a need to develop novel approaches that are more robust for unseen environments.
@@ I-B Contributions
    Extensive simulations demonstrate that the proposed SCA-LLM achieves state-of-the-art (SOTA) prediction performance and exhibits strong robustness and generalization capabilities, across diverse communication scenarios, mobility conditions, and signal-to-noise ratios (SNRs). Ablation studies further validate the effectiveness and superiority of the proposed SCA-LLM, confirming its significant contribution to adapting the LLM for channel prediction task.
@@ II-A MIMO-OFDM System
In this work, we consider a MIMO-OFDM communication system where a base station (BS) equipped with $N_{BS}$ antennas communicates with user equipments (UEs) with $N_{UE}$ antennas. In order to optimize the data transmission to the users, the BS must conduct CSI estimation for all users. Assuming orthogonal pilot sequences across the users, the CSI estimation and prediction procedure can be applied to every user independently. Without loss of generality, we focus on a particular active user in the system, thus omit the user index [20]. The total bandwidth $B$ is divided into $K$ orthogonal subcarriers, indexed by $k=0,1,\dots,K-1$, with subcarrier spacing $\Delta f$. A cyclic prefix (CP), configured to be longer than the maximum channel delay spread, is added to each OFDM symbol. Moreover, the system operates in time-division duplexing (TDD) mode, where the uplink (UL) and downlink (DL) transmissions use the same frequency band but occur at different time slots within a communication frame. Each frame allocates specific time slots for transmitting uplink pilot signals, such as sounding reference signals (SRS) from the UE. These pilots enable the BS to periodically estimate the UL channel state information (CSI). Following the pilot transmissions, the frame structure provides configurable time slots designated for UL and DL data transmission.
@@ II-B Channel Model
$\begin{split}\mathbf{H}_{UL}(t,f)={}&\sum_{n=1}^{N}\sum_{m=1}^{M}\alpha_{n,m}e^{-j2\pi f\tau_{n,m}}e^{j2\pi\nu_{n,m}t}\\ &\quad\mathbf{a}_{BS}(\theta_{n,m,ZOA},\phi_{n,m,AOA})\\ &\qquad\mathbf{a}_{UE}^{H}(\theta_{n,m,ZOD},\phi_{n,m,AOD}),\end{split}$ (1)
@@ V Numerical Results
This section describes the environment setup, baseline methods, training details, evaluation results, and ablation studies. The source code of this paper can be found at https://github.com/hele082/SCA-LLM.
@@ V-D1 NMSE Performance Comparison under Varying SNR
(a) UMa NLOS Scenario
(b) UMi NLOS Scenario
@@
(a) UMa NLOS Scenario
(b) UMi NLOS Scenario
@@
Focusing first on the static scenario ($0~\text{km/h}$), which is crucial for applications like fixed wireless access or low-mobility IoT devices, Fig. 2 details the performance comparison. In the UMa NLOS scenario (Fig. 2a), corresponding to the training distribution, our proposed method significantly outperforms all baselines. For instance, at an SNR of $20~\text{dB}$, our method demonstrates a performance advantage of approximately $2.0~\text{dB}$ over the next best baseline, LLM4CP (achieving NMSEs of roughly $22.3~\text{dB}$ versus $20.3~\text{dB}$, respectively), and is substantially better than the traditional methods (RNN, LSTM, GRU, Transformer). When evaluating the generalization capability in the unseen UMi NLOS scenario (Fig. 2b), our method maintains its clear performance leadership. Interestingly, while all methods tend to exhibit a slight improvement in NMSE performance when generalizing from the trained UMa environment to the unseen UMi environment, our method still consistently achieves the lowest NMSE across the tested SNR range. Furthermore, the performance gap between our method and LLM4CP widens in the zero-shot UMi setting. In particular, at $20~\text{dB}$ SNR, this gap increases from approximately $2.0~\text{dB}$ in UMa to $2.4~\text{dB}$ in UMi, where the respective NMSEs are $22.7~\text{dB}$ (our method) and $20.3~\text{dB}$ (LLM4CP). This further highlights the superior robustness and generalization of our proposed framework.
@@ V-D2 NMSE Performance Comparison under Varying UE Velocity
(a) UMa NLOS Scenario
(b) UMi NLOS Scenario
@@ V-D3 NMSE Performance per Prediction Step Comparison
(a) UMa NLOS Scenario
(b) UMi NLOS Scenario
@@ VI Conclusions
This paper proposed the SCA-LLM framework for accurate and robust channel prediction in MIMO-OFDM systems. The framework effectively adapted pre-trained LLMs to CSI data by mitigating domain mismatch through a novel spectral-attentive adapter. In this adapter, the SCA module employed multi-spectral channel attention with the 2-D DCT to capture a broader range of spectral components from CSI features compared to methods relying on global pooling operations. This, in turn, provided a more suitable input for the LLM and significantly improved its adaptation and prediction capabilities. Extensive simulations demonstrated that SCA-LLM achieved superior prediction performance, consistently outperformed existing methods, and yielded up to a notable NMSE advantage of $2.4~\text{dB}$ over the prior LLM-based approach. Ablation studies further validated the critical contributions of both the spectral-attentive adapter and the LLM backbone. These findings underscored that the SCA-LLM framework strengthened the “adapter + LLM” paradigm, enabled better exploitation of LLMs for channel prediction, and suggested potential for broader LLM applications in wireless communications.

### v2

# SCA-LLM: Spectral-Attentive LLM-Based Wireless World Modeling for Agentic Communications

@@ Abstract
Future AI-native wireless networks are moving from reactive optimization to agentic decision-making that can sense, predict, and plan under fast-varying channels. This calls for wireless world models that can predict and roll out channel dynamics, for which multi-step channel state information (CSI) prediction offers a practical short-horizon look-ahead. Recent advances in foundation sequence models further motivate large language models (LLMs) as general-purpose dynamics learners when suitably adapted to non-text time-series signals. However, bridging CSI to LLMs is non-trivial because an effective adapter must expose informative spectral and temporal evolution patterns, while prior designs provide limited inductive bias to capture such channel structures. To this end, we propose SCA-LLM, a spectral-attentive LLM-based wireless world modeling framework that bridges CSI to LLMs via a spectral-channel attention (SCA) adapter. Specifically, the SCA adapter performs multi-spectral representation learning to extract informative channel features and align CSI with the LLM’s sequence modeling capability, enabling parameter-efficient adaptation while keeping the LLM backbone largely frozen. Extensive simulations show that SCA-LLM achieves state-of-the-art prediction performance and strong zero-shot generalization, yielding up to $2.4~\text{dB}$ normalized mean squared error (NMSE) advantage over the previous LLM based method. Our ablation studies further confirm the effectiveness of the proposed SCA adapter in mitigating domain mismatch.
@@ I Introduction
Artificial intelligence-radio access network (AI-RAN) is transforming wireless system design from hand-crafted model-driven optimization toward data-driven agentic communications [1, 2]. In this paradigm, the network is expected to behave as an agentic system that can continuously sense, predict, and perform look-ahead planning under fast-varying propagation conditions, rather than merely reacting to the instantaneous channel state [3, 4]. For modern multiple-input multiple-output orthogonal frequency division multiplexing (MIMO-OFDM) deployments [5, 6, 7], adaptive techniques such as dynamic resource allocation [8], antenna selection [3], precoding [9], and beamforming [10] are increasingly time-sensitive and uncertainty-aware. However, practical CSI acquisition and feedback inevitably incur latency and overhead [4], and the resulting channel aging makes instantaneous CSI insufficient to reliably guide uncertainty-aware optimizations in dynamic environments [11]. Therefore, the ability to forecast the multi-step evolution of channel state information (CSI) becomes a key enabler for reliable and efficient look-ahead decision-making in agentic communications [3, 12].
From this perspective, multi-step CSI prediction naturally serves as a short-horizon wireless world model that predicts the transition dynamics of the propagation environment. By rolling out future CSI trajectories, the network agent gains the ability to reason about the near-future channel evolution and to proactively plan transmission actions, rather than reacting to a possibly stale CSI snapshot [3]. Recognizing this, recent research has begun to leverage wireless world models for agentic communications. For example, the authors in [3] and [4] proposed to use deep generative models as the wireless world model, and perform risk-aware antenna selection via look-ahead planning on the near-future CSI evolutions.
Despite these encouraging results, existing wireless world models are typically small-scale and built upon scenario-specific models [3, 4, 13, 14, 15, 16, 17, 18, 19, 20], whose generalization is often limited when facing heterogeneous deployments, varying mobility patterns, and diverse radio configurations. As AI-RAN scales up and operates in continually changing environments, we increasingly need a reusable foundational sequence-model backbone for wireless world modeling, which should encode strong dynamics priors, transfer well across scenarios, and admit parameter-efficient adaptation. This is also the main focus of this work.
@@ I-A Prior Works
Existing channel prediction methods can be broadly categorized into traditional methods and deep learning (DL) based methods. Traditional methods are typically based on explicit mathematical models derived from statistical assumptions or simplified representations of the underlying channel physics. However, it is often difficult for these methods to perform well in complex and dynamic wireless environments. For instance, although advanced model-based techniques, such as the Prony-based angle-delay domain (PAD) method [13], are designed to capture specific channel structures and potentially offer improved accuracy under certain conditions, fundamental limitations are often encountered. While analytical guarantees are provided by PAD, frequent updates to its model parameters are required for adaptation to environmental changes[14]. Substantial computational overhead is incurred by this process as well. There is still a need for more robust and generalizable channel prediction approaches capable of effective adaptation to diverse environments.
@@
In response to this need, DL based methods [14, 15, 16, 17, 18, 19, 20] have emerged as a promising alternative. Rather than relying on predefined mathematical models, DL methods are data-driven approaches that learn complex dependencies and predictive patterns directly from historical channel observations. Architectures such as recurrent neural networks (RNNs) [16], long short-term memory (LSTM) networks [18], gated recurrent units (GRUs) [19], and transformers [20] have been employed, attempting to model the complex spatio-temporal dynamics of wireless channels. For example, focusing on the angle-delay domain for high-mobility massive MIMO systems, a complex-valued neural network (CVNN) based prediction method has been introduced in [14]. This data-driven approach performs element-wise prediction of the angle-delay channel response matrix (ADCRM). By learning directly from historical channel data represented in the angle-delay domain, this method achieved improved prediction accuracy compared to the traditional model-based PAD method [13]. Furthermore, transformer-based methods for channel prediction have been introduced in [4] and [20]. These approaches use the attention mechanism within the transformer model to establish the relationship between past CSI and future channels. It gives more importance to the most relevant past information, which improves the prediction accuracy and overall communication performance. Additionally, a unified pilot-to-precoder scheme has been proposed either in [20], using a single transformer model to integrate channel estimation, prediction, and precoding, thereby reducing computational complexity. While these approaches have often demonstrated improved accuracy over traditional methods, they still have limitations in practice. Specifically, these models are typically trained for specific tasks and environments. When encountering unseen communication environments without significant retraining, they often exhibit a lack of robustness due to limited sequence modeling capabilities. Therefore, there is a need to develop novel approaches that are more robust for unseen environments.
@@ I-B Contributions
    Extensive simulations demonstrate that the proposed SCA-LLM achieves state-of-the-art (SOTA) prediction performance and exhibits strong robustness and zero-shot generalization capabilities, across diverse communication scenarios, mobility conditions, and signal-to-noise ratios (SNRs). Ablation studies further validate the effectiveness and superiority of the proposed SCA-LLM, confirming its significant contribution to adapting the LLM for channel prediction task.
@@ II-A MIMO-OFDM System
In this work, we consider a MIMO-OFDM communication system where a base station (BS) equipped with $N_{BS}$ antennas communicates with user equipments (UEs) with $N_{UE}$ antennas. In order to optimize the data transmission to the users, the BS must conduct CSI estimation for all users. Assuming orthogonal pilot sequences across the users, the CSI estimation and prediction procedure can be applied to every user independently. Without loss of generality, we focus on a particular active user in the system, thus omit the user index [20, 33]. The total bandwidth $B$ is divided into $K$ orthogonal subcarriers, indexed by $k=0,1,\dots,K-1$, with subcarrier spacing $\Delta f$. A cyclic prefix (CP), configured to be longer than the maximum channel delay spread, is added to each OFDM symbol. Moreover, the system operates in time-division duplexing (TDD) mode, where the uplink (UL) and downlink (DL) transmissions use the same frequency band but occur at different time slots within a communication frame. Each frame allocates specific time slots for transmitting uplink pilot signals, such as sounding reference signals (SRS) from the UE. These pilots enable the BS to periodically estimate the UL channel state information (CSI). Following the pilot transmissions, the frame structure provides configurable time slots designated for UL and DL data transmission.
@@ II-B Channel Model
$\begin{split}\mathbf{H}_{UL}(t,f)={}&\sum_{n=1}^{N}\sum_{m=1}^{M}\alpha_{n,m}e^{-j2\pi f\tau_{n,m}}e^{j2\pi\nu_{n,m}t}\mathbf{a}_{BS}(\theta_{n,m,ZOA},\phi_{n,m,AOA})\mathbf{a}_{UE}^{H}(\theta_{n,m,ZOD},\phi_{n,m,AOD}),\end{split}$ (1)
@@ V Numerical Results
This section describes the environment setup, baseline methods, training details, evaluation results, and ablation studies  [^1] Our source code is available at https://github.com/hele082/SCA-LLM..
@@ V-D1 NMSE Performance Comparison under Varying SNR
(a) UMa NLOS scenario (trained from scratch)
(b) UMi NLOS scenario (zero-shot evaluation; trained on UMa)
@@
(a) UMa NLOS scenario (trained from scratch)
(b) UMi NLOS scenario (zero-shot evaluation; trained on UMa)
@@
Focusing first on the static scenario ($0~\text{km/h}$), which is crucial for applications like fixed wireless access or low-mobility IoT devices, Fig. 2 details the performance comparison. In the UMa NLOS scenario (Fig. 2a), corresponding to the training distribution, our proposed method significantly outperforms all baselines. For instance, at an SNR of $20~\text{dB}$, our method demonstrates a performance advantage of approximately $2.0~\text{dB}$ over the next best baseline, LLM4CP (achieving NMSEs of roughly $22.3~\text{dB}$ versus $20.3~\text{dB}$, respectively), and is substantially better than the traditional methods (RNN, LSTM, GRU, Transformer).
When evaluating the generalization capability in the unseen UMi NLOS scenario (Fig. 2b), our method maintains its clear performance leadership. Interestingly, while all methods tend to exhibit degradation in NMSE performance when generalizing from the trained UMa environment to the unseen UMi environment, our method still consistently achieves the lowest NMSE across the tested SNR range. Furthermore, the performance gap between our method and LLM4CP widens in the zero-shot UMi setting. In particular, at $20~\text{dB}$ SNR, this gap increases from approximately $2.0~\text{dB}$ in UMa to $2.4~\text{dB}$ in UMi, where the respective NMSEs are $22.7~\text{dB}$ (our method) and $20.3~\text{dB}$ (LLM4CP). This further highlights the superior robustness and generalization of our proposed framework.
@@ V-D2 NMSE Performance Comparison under Varying UE Velocity
(a) UMa NLOS scenario (trained from scratch)
(b) UMi NLOS scenario (zero-shot evaluation; trained on UMa)
@@ V-D3 NMSE Performance per Prediction Step Comparison
(a) UMa NLOS scenario (trained from scratch)
(b) UMi NLOS scenario (zero-shot evaluation; trained on UMa)
@@ VI Conclusions
This paper proposed the SCA-LLM framework, which is a large-scale wireless world model for accurate and robust multi-step CSI prediction in MIMO-OFDM systems. The framework effectively adapted pre-trained LLMs to CSI data by mitigating domain mismatch through a novel spectral-attentive adapter. In this adapter, the SCA module employed multi-spectral channel attention with the 2-D DCT to capture a broader range of spectral components from CSI features compared to methods relying on global pooling operations. This, in turn, provided a more suitable input for the LLM and significantly improved its adaptation and prediction capabilities. Extensive simulations demonstrated that SCA-LLM achieved superior prediction performance, consistently outperformed existing methods, and yielded up to a notable NMSE advantage of $2.4~\text{dB}$ over the prior LLM-based approach. Ablation studies further validated the critical contributions of both the spectral-attentive adapter and the LLM backbone. These findings underscored that the SCA-LLM framework strengthened the “adapter + LLM” paradigm, enabled better exploitation of LLMs for channel prediction, and suggested potential for broader LLM applications in wireless communications.